OpenStudy (anonymous):

x^2-2/3x (same instructions) @satellite73

OpenStudy (anonymous):

the fractions kind of scaring me

OpenStudy (anonymous):

half of \(\frac{2}{3}\) is ?

OpenStudy (anonymous):

rude

OpenStudy (anonymous):

don't let them scare you just take half of two thirds

OpenStudy (anonymous):

hm 1/3?

OpenStudy (anonymous):

not 1/3? but rather 1/3!!

OpenStudy (anonymous):

that is step one step two is what is \(\left(\frac{1}{3}\right)^2\)

OpenStudy (anonymous):

okay so 1/9

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so \[x^2-\frac{2}{3}x+\frac{1}{9}=(x-\frac{1}{3})^2\]

OpenStudy (anonymous):

these seem a lot simpler than I thought. Thanks for helping me realize how to find the answer and how to not overthink things and doubt

OpenStudy (anonymous):

they are very simple once you know the rules really easy even with (ugh) fractions

OpenStudy (anonymous):

could we do one more?

OpenStudy (anonymous):

ok why not?

OpenStudy (anonymous):

so x^2-1/3x

OpenStudy (anonymous):

ooh another fraction want to try it?

OpenStudy (anonymous):

half of 1/is 1/6 right? 1/3 x 1/2? (so much for not doubting)

OpenStudy (anonymous):

right, half of \(\frac{1}{3}\) is \(\frac{1}{6}\)

OpenStudy (anonymous):

then second , what is \((\frac{1}{6})^2\) ?

OpenStudy (anonymous):

x^2-1/3+1/12= (x^2-1/6)2

OpenStudy (anonymous):

oops so close

OpenStudy (anonymous):

you balked on \[\huge \left(\frac{1}{6}\right)^2\]

OpenStudy (anonymous):

x^2 -1/3x+1/36=(x^2-1/6)^2 WOW..

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks. Shall we complete the square now? I just need some help getting started

OpenStudy (anonymous):

lol what the heck you think you have been doing???

OpenStudy (anonymous):

that's what the second set of instructions is asking for specifically so I am going by what my textbook says

OpenStudy (anonymous):

ok lets try one believe me it is the same

OpenStudy (anonymous):

okay let's do this x^2+6x=7

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

you would still take hald of six and square it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

*half

OpenStudy (anonymous):

subtract 7 over?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

leave it right where it is

OpenStudy (anonymous):

you want me to walk you though this one?

OpenStudy (anonymous):

x^2+6x+9=7?

OpenStudy (anonymous):

ok you started with \[x^2+6x=7\] when you complete the square as you have been doing you get \[x^2+6x+9\] but you have to add \(9\) to the right as well so it is \[x^2+6x+9=7+9\\ (x+3)^2=16\]

OpenStudy (anonymous):

then it is a snap to solve \[(x+3)^2=16\\x+3=4,x+3=-4\\ x=1,x=-7\]

OpenStudy (anonymous):

i see said the blind man

OpenStudy (anonymous):

to his deaf wife over the disconnected telephone as he stuck his head out of the closed window to see if it was raining

OpenStudy (anonymous):

you can really go right from \[x^2+6x=7\] to \[(x+3)^2=7+9\] if you like

OpenStudy (anonymous):

i am tired want to try one more?

OpenStudy (anonymous):

so for x^2-2x=2 x^2-2x+1=2+1 (x^2+2)^2=3? I think I just confused myself

OpenStudy (anonymous):

everything good but the last line

OpenStudy (anonymous):

\[(x-1)^2=3\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[x^2-2x=2\\ x^2-2x+1=2+1\\ (x-1)^2=3\]

OpenStudy (anonymous):

just like \[x^2+6x=7\\ x^2+6x+9=7+9\\ (x+3)^2=16\]

OpenStudy (anonymous):

just foiled out the (x-1)^2 part , I don't know why but it threw me off. Anyway, I don't know how else to thank you, you've been very helpful. Also, sorry for hogging you @satellite73 goodnight!

OpenStudy (anonymous):

hey no problem gnight to you